Richie put on his jogging shoes and speed-walked to the park at 5 miles per hour. Then he ran back at 9 miles per hour. If the total trip took 4 2/3 hours, how far was it to the park?
Answer by checkley75(3666)

D/5+D/9=4 2/3
(9D+5D)/45=14/3 NOW CROSS MULTIPLY
14D*3=14*45
42D=630
D=630/42
D=15 MILES IS THE DISTANCE TO THE PARK.
PROOF:
15/5+15/9=4 2/3
3+1 2/3=4 2/3
4 2/3=4 2/3